# Language-free Compositional Action Generation via Decoupling Refinement

### v1

@@ 4 Experiments
4.1 Dataset & Evaluation Metrics
@@
HumanAct-C dataset: The HumanAct-C dataset is organized into a training set of original sub-actions and a testing set of compositional actions. The training set is derived from a subset of the original HumanAct12 training set. As for the testing set, we synthesize compositional actions using testing samples from a pair of action sequences with differing labels. These sequences are fused using motion energy-based attention masks, and manually curated to remove any irrational samples and unfeasible combination categories, such as “drinking while phoning”. In the end, we curate a training set of 120 sequences for 12 sub-actions, and a testing set featuring 9 compositional categories with a total of 900 compositional action sequences. Please note that the sub-action categories in training and testing sets may not be strictly aligned. For example, there are 12 sub-actions for training, and 7 of them are used to build compositional actions.
UESTC-C dataset: We adopt a similar approach to construct the UESTC-C dataset. We generate the testing set comprising 25 compositional categories with a total of $25\times 25=625$ compositional action sequences. The training set is composed of $10\times 5=50$ sequences. Both the training and testing sets consist of the same 10 sub-actions. More visualizations and detailed descriptions of these datasets are available in the supplementary materials.
@@ 4.2 Implementation Details
The Transformer decoder utilizes a combination of the latent vector and the compositional label embedding for its key and value inputs, while a sequence of positional embeddings is used as the query. The decoder’s output is a sequence of 3D actions, matching the size of the input pseudo-actions. To further refine these 3D actions and acquire the corresponding body surfaces, we can employ the SMPL model.
Hyper-parameters:
The final loss function of our model comprises three components: the reconstruction loss and the KL loss as per Eq. (5), and the Decoupling Refinement (DR) loss as defined by Eq. (7). The balance rate between these components is set as $21:1e-5:1e-$ 2. To ensure an efficient training, we follow ACTOR’s [23] example by setting the frame length to 64, with each frame containing $N=24$ joint points. These points are projected into a continuous 6D rotation representation space $\mathbb{R}^{24\times 6}$. During the action coupling process, the standard deviation $\sigma$ of the Gaussian distribution is set to 0.1. Both the encoder and the decoder employ eight transformer layers. Regarding the MAE model, we utilize a larger vision that also incorporates a GAN loss to enhance the in-painting capabilities.
@@
4.3 Quantitative Evaluation
We compare our method with three main baseline methods including Pre-train+Mask [23], Latent Disentanglement [1], and the text-guided method MotionCLIP [30]. Pre-train+Mask is an intuitive solution for the compositional generation to directly compose two generated actions with attention masks. Latent disentanglement learns the latent variables for more diverse generations. However, they didn’t involve the action coupling process to produce more diverse training data and the decoupling refinement process to improve the action quality. The text-guided methods have been discussed in Section 2. Please refer to the supplementary materials for more details on the baseline methods.
@@ 4.3 Quantitative Evaluation
Figure 3: Compositional 3D motion generations with 9 categories on HumanAct-C. We provide the visualizations of a pair of raw sub-actions and the compositional actions generated by our method. Zoom in for details.
4.4 Ablation Studies
We also conduct ablation studies to investigate the effectiveness of different components. The performance results are summarized in Table 2, which serves as the evidence to answer the following questions:
Q: Can the pre-trained text-guided methods be successfully applied to our scenarios? A: No. As indicated in Table 1, MotionCLIP [30] struggles to generate compositional actions when its training data lacks corresponding compositional actions. Specifically, in our experiments, we utilize the MotionCLIP model pre-trained on AMASS [20] (we’re unable to train MotionCLIP using our dataset due to the absence of language annotations and compositional data). During inference, we use the phrase "action1 while action2" as the text prompt for simultaneous compositional actions for MotionCLIP. However, the results show that the model fails to generate satisfactory results when the test data extends beyond the scope of the training text. This is further evidenced by the visualizations provided in Figure 6.
Figure 4: Compositional 3D motion generations with different categories on the UESTC-C dataset. We only display 5 compositional categories here and put the full visualization in the supplementary materials. Please zoom in for the details.
Q: What is the effect of the energy-based attention mask? A: It is the key to coupling and decoupling. In an experimental setting where the attention mask was set to $\bm{1}$, indicating equal scores for all body parts, the performance significantly declined, with an FID score of 251.85 and an accuracy rate of 31.6. This was notably worse than baseline methods. This finding underscores the importance of the attention mask for learning disentangled representations, as crucial attributes of sub-actions tend to interact with each other. Additionally, a visual comparison, which further illustrates the impact of our energy-based attention mask, has been provided in the supplementary materials.
Figure 5: Visualization using SMPL with different sampled sequences of the same compositional categories.
4.5 Qualitative Evaluation
In this subsection, we showcase several visualization examples of the generated compositional 3D actions. Figure 3 displays results for nine categories from the HumanAct-C dataset. For instance, observing the "drink and walk" example, we see the generation of realistic compositional 3D motions. This validates our method’s efficacy in learning the disentangled latent variable and its applicability to zero-shot compositional action generation. Moreover, Figure 5 presents multiple sequences for a single compositional category using SMPL [18] to highlight the diversity of our generations, where two sequences each for "walk and drink" and "sit and drink" are displayed. Additionally, we offer some visualization examples of compositional 3D actions generated on the UESTC-C datasets. Figure 4 exhibits results for 5 out of the 25 categories from the UESTC-C datasets. A comprehensive visualization of all 25 categories is available in the supplementary materials. Figure 6 compares the generations among the Pre-trained ACTOR [23] + Mask, MotionCLIP [30], and ours.
Figure 6: The qualitative comparisons between our methods and other baseline methods. Discussions are provided in the appendix.
5 Conclusion
In this paper, we have presented an innovative framework for the generation of compositional 3D action sequences, without the need for extensive annotations. Our framework comprises three primary components: Action Coupling, Conditional Action Generation, and Decoupling Refinement. Through the use of an energy model, we extract attention masks and create pseudo-compositional actions, and then employ a conditional generative model to learn a diverse latent space. Furthermore, Decoupling Refinement ensures semantic consistency between the sub-actions and compositional actions. We have also built two novel datasets, HumanAct-C and UESTC-C, and corresponding evaluation metrics, providing a valuable resource for future research in this domain. Comprehensive evaluations of these datasets have shown our methodology’s significant effectiveness in generating compositional actions, opening new avenues for further research in the field of 3D action generation.
Limitation & Border Impact: Since this is the original attempt for 3D compositional action generation without text annotations, we conduct experiments on datasets with limited categories. In the future, we will extend the scale of the constructed benchmarks. For the potential negative societal impacts, people should avoid adopting this method to generate fake data for unlawful use.
APPENDIX
A.1 More Implementation Details
In the presented experiments, an 8-layer Encoder and Decoder were utilized within the Transformer model. Multi-head attention was configured with 4 heads, and a dropout rate of 0.1 was applied to prevent overfitting and enhance the generalization ability of the model. The input embedding dimension was set at 256, and the intermediate feedforward network had a dimension of 1024.
Figure A.1: The visualization of examples of our dataset HumanAct-C and UESTC-C.
A.2 Datasets
HumanAct-C dataset: The original HumanAct12 dataset includes 12 action categories, annotated with joint coordinates and SMPL pose parameters processed by [23]. We utilized these categories as sub-actions, from which we can generate compositional action categories by combining each pair of actions to build the HumanAct-C dataset. HumanAct-C is organized into a training set of original sub-actions and a testing set of compositional actions. The training set is derived from a subset of the original HumanAct12 training set. As for the testing set, we synthesize compositional actions using testing samples from a pair of action sequences with differing labels. These sequences are fused using motion energy-based attention masks, and manually curated to remove any irrational samples and unfeasible combination categories, such as “drinking while phoning”. In the end, we curate a training set of 120 sequences for 12 sub-actions, and a testing set featuring 9 compositional categories with a total of 900 compositional action sequences. These 9 compositional categories are based on 7 commonly seen sub-actions: "walk+drink", "walk+call", "walk+throw", "walk+ lift dumbbell", "run+call", "run+throw", "sit+call", "sit+eat", and "sit+drink". Please note that the sub-action categories in training and testing sets may not be strictly aligned. For example, there are 12 sub-actions for training, and 7 of them are used to build compositional actions. The examples of these 7 compositional actions are visualized in Figure A.1.
Figure A.2: Compositional 3D motion generations with different categories. We provide the visualizations of a pair of raw sub-actions and our generated compositional actions on all 25 compositional categories. Zoom in for the details.
Figure A.3: The qualitative comparisons between our methods and other baseline methods.
A.3 Comprehensive Ablation Study Results
In Table A.1, we display the comprehensive outcomes of our ablation studies, which further include a meticulous examination of our model’s performance based on the diversity metric. The detailed results attest to the efficacy of our method, demonstrating that it accomplishes more accurate generation without compromising the diversity of the results.
@@ A.3 Comprehensive Ablation Study Results
A.4 Complete Qualitative Results on UESTC-C
We provide comprehensive visualizations of our model’s performance on the UESTC-C dataset in Figure A.2. The dataset comprises all 25 categories of actions. The visual results substantiate our approach’s effectiveness, demonstrating our model’s capability to generate clear and reasonable composite actions using solely sub-actions as training inputs. This observation further attests to the flexibility and adaptability of our model when addressing complex tasks.
A.5 Comparisons with Baseline Methods
We compare our method with three main baseline methods including Pre-train+Mask [23], Latent Disentanglement [1], and the text-guided method MotionCLIP [30]. The quantitative comparisons can be found in Section 4.3 in the main submission. Below, we show more details of the implementation of baseline methods and qualitative comparisons.
We juxtapose our method with three primary baseline methods: Pre-train+Mask (ACTOR)[23], Latent Disentanglement[1], and the text-guided method MotionCLIP [30]. Quantitative comparisons can be found in Section 4.3 of the main submission. In this section, we provide additional details regarding the implementation of these baseline methods and offer qualitative comparisons.
@@ A.5 Comparisons with Baseline Methods
Detailed comparisons with these baseline methods are provided in Section 4.3 and Figure 6 in the main manuscript. For convenience, we reproduce the same figure here as Figure A.3 and provide additional explanations regarding these qualitative results. Firstly, we noticed that MotionCLIP does not generate actions that are easily distinguishable or adequately representative of the intended behaviors. For instance, the actions corresponding to "sitting + calling" and "sitting + drinking" are quite similar with limited variations, making it difficult to identify the distinct characteristics of calling or drinking in the animations. On the other hand, ACTOR+Mask manages to generate reasonable composite actions, but these actions lack fluidity and some characteristic elements are not clearly defined. Taking "walking + drinking" as an example, the animation does not exhibit a clear walking motion. By contrast, our proposed method demonstrates an improved capacity to generate more discernible and smooth composite actions, highlighting the key features of each individual action. This comparison underscores the superiority of our approach in generating coherent and representative composite actions. Please note that we use a white background specifically for visualizations involving MotionCLIP to indicate that it was not trained on our dataset. However, it’s important to emphasize that the choice of the background color is purely cosmetic and does not have any bearing on the quantitative performance of the model.
@@
Figure A.4: The qualitative comparisons between our methods and the one without motion energy for building attention.
A.6 Qualitative Evaluation of Motion Energy
@@
In Section 4.4 and Table 2, we conducted a quantitative analysis of the influence of motion energy. In this addendum, we supplement this analysis with visual comparisons displayed in Figure A.4. A cursory visual examination reveals that models without the application of motion energy may fail to capture the crucial characteristics of the sub-actions. For instance, in the "run + call" action, the model neglects the characteristic motion of making a phone call when motion energy is not used to emphasize the significant parts of the sub-action. This further demonstrates the critical role that motion energy plays in accurately capturing and preserving the distinctive characteristics of each sub-action in the final composed motion.

### v2

@@ A.3 More Experiments
Fig. A.1: The visualization of examples of our dataset HumanAct-C and UESTC-C.
A.3.1 Dataset & Evaluation Metrics
@@
HumanAct-C dataset: The original HumanAct12 dataset includes 12 action categories, annotated with joint coordinates and SMPL pose parameters processed by [2]. We utilized these categories as sub-actions, from which we can generate compositional action categories by combining each pair of actions to build the HumanAct-C dataset. HumanAct-C is organized into a training set of original sub-actions and a testing set of compositional actions. The training set is derived from a subset of the original HumanAct12 training set. As for the testing set, we synthesize compositional actions using testing samples from a pair of action sequences with differing labels. These sequences are fused using motion energy-based attention masks, and manually curated to remove any irrational samples and unfeasible combination categories, such as “drinking while phoning”. In the end, we curate a training set of 120 sequences for 12 sub-actions, and a testing set featuring 9 compositional categories with a total of 900 compositional action sequences. These 9 compositional categories are based on 7 commonly seen sub-actions: "walk+drink", "walk+call", "walk+throw", "walk+ lift dumbbell", "run+call", "run+throw", "sit+call", "sit+eat", and "sit+drink". Please note that the sub-action categories in training and testing sets may not be strictly aligned. For example, there are 12 sub-actions for training, and 7 of them are used to build compositional actions. The examples of these 9 compositional actions are visualized in Figure A.1.
@@ A.3.2 Implementation Details
The Transformer decoder utilizes a combination of the latent vector and the compositional label embedding for its key and value inputs, while a sequence of positional embeddings is used as the query. The decoder’s output is a sequence of 3D actions, matching the size of the input pseudo-actions. To further refine these 3D actions and acquire the corresponding body surfaces, we employ the SMPL model.
Hyper-parameters: The final loss function of our model comprises three components: the reconstruction loss and the KL loss as per Eq. (12), and the Decoupling Refinement (DR) loss as defined by Eq. (14). The balance rate between these components is set as $21:1e-5:1e-$ 2. To ensure an efficient training, we follow ACTOR’s [2] example by setting the frame length to 64, with each frame containing $N=24$ joint points. These points are projected into a continuous 6D rotation representation space $\mathbb{R}^{24\times 6}$. During the action coupling process, the standard deviation $\sigma$ of the Gaussian distribution is set to 0.1. In the presented experiments, an 8-layer Encoder and Decoder were utilized within the Transformer model. Multi-head attention was configured with 4 heads, and a dropout rate of 0.1 was applied to prevent overfitting and enhance the generalization ability of the model. The input embedding dimension was set at 256, and the intermediate feedforward network had a dimension of 1024. Regarding the MAE model, we utilize a larger vision that also incorporates a GAN loss to enhance the in-painting capabilities.
@@
A.3.3 Quantitative Evaluation
We compare our method with three main baseline methods including Pre-train+Mask (ACTOR)[2], Latent Disentanglement [4], and the text-guided method MotionCLIP [5]. Pre-train+Mask is an intuitive solution for the compositional generation to directly compose two generated actions with attention masks. Latent disentanglement learns the latent variables for more diverse generations. However, they didn’t involve the action coupling process to produce more diverse training data and the decoupling refinement process to improve the action quality. The text-guided methods have been discussed in Section A.1.
@@ A.3.3 Quantitative Evaluation
Fig. A.2: Compositional 3D motion generations with 9 categories on HumanAct-C. We provide the visualizations of a pair of raw sub-actions and the compositional actions generated by our method. Zoom in for details.
A.3.4 Ablation Studies
@@
We also conduct ablation studies to investigate the effectiveness of different components. In Table A.1, we display the comprehensive outcomes of our ablation studies, which further include a meticulous examination of our model’s performance based on the diversity metric. The detailed results attest to the efficacy of our method, which serves as the evidence to answer the following questions:
Q: Can the pre-trained text-guided methods be successfully applied to our scenarios? A: No. As indicated in Table 1, MotionCLIP [5] struggles to generate compositional actions when its training data lacks corresponding compositional actions. Specifically, in our experiments, we utilize the MotionCLIP model pre-trained on AMASS [25] (we’re unable to train MotionCLIP using our dataset due to the absence of language annotations and compositional data). During inference, we use the phrase "action1 while action2" as the text prompt for simultaneous compositional actions for MotionCLIP. However, the results show that the model fails to generate satisfactory results when the test data extends beyond the scope of the training text. This is further evidenced by the visualizations provided in Figure A.4.
Q: What is the effect of the energy-based attention mask? A: It is the key to coupling and decoupling. In an experimental setting where the attention mask was set to $\bm{1}$, indicating equal scores for all body parts, the performance significantly declined, with an FID score of 251.85 and an accuracy rate of 31.6. This was notably worse than baseline methods. This finding underscores the importance of the attention mask for learning disentangled representations, as crucial attributes of sub-actions tend to interact with each other. Additionally, a visual comparison, which further illustrates the impact of our energy-based attention mask, has been provided in Figure A.5.
Fig. A.3: Visualization using SMPL with different sampled sequences of the same compositional categories.
A.3.5 Qualitative Evaluation
In this subsection, we showcase several visualization examples of the generated compositional 3D actions. Figure A.2 displays results for nine categories from the HumanAct-C dataset. For instance, observing the "drink and walk" example, we see the generation of realistic compositional 3D motions. This validates our method’s efficacy in learning the disentangled latent variable and its applicability to zero-shot compositional action generation. Moreover, Figure A.3 presents multiple sequences for a single compositional category using SMPL [24] to highlight the diversity of our generations, where two sequences each for "walk and drink" and "sit and drink" are displayed. Additionally, We provide comprehensive visualizations of our model’s performance on the UESTC-C dataset in Figure A.6. The dataset comprises all 25 categories of actions.
@@ A.3.5 Qualitative Evaluation
Figure A.4 compares the generations among the Pre-trained ACTOR [2] + Mask, MotionCLIP [5], and ours. We notice that MotionCLIP does not generate actions that are easily distinguishable or adequately representative of the intended behaviors. For instance, the actions corresponding to "sitting + calling" and "sitting + drinking" are quite similar with limited variations, making it difficult to identify the distinct characteristics of calling or drinking in the animations. On the other hand, ACTOR+Mask manages to generate reasonable composite actions, but these actions lack fluidity and some characteristic elements are not clearly defined. Taking "walking + drinking" as an example, the animation does not exhibit a clear walking motion. By contrast, our proposed method demonstrates an improved capacity to generate more discernible and smooth composite actions, highlighting the key features of each individual action. This comparison underscores the superiority of our approach in generating coherent and representative composite actions. Please note that we use a white background specifically for visualizations involving MotionCLIP to indicate that it was not trained on our dataset. However, it’s important to emphasize that the choice of the background color is purely cosmetic and does not have any bearing on the quantitative performance of the model.
Fig. A.4: The qualitative comparisons between our methods and other baseline methods.
@@
Fig. A.5: The qualitative comparisons between our methods and the one without motion energy for building attention.
Fig. A.6: Compositional 3D motion generations with different categories. We provide the visualizations of a pair of raw sub-actions and our generated compositional actions on all 25 compositional categories. Zoom in for the details.
@@
In Section A.3.4 and Table A.1, we conducted a quantitative analysis of the influence of motion energy. In this addendum, we supplement this analysis with visual comparisons displayed in Figure A.5. A cursory visual examination reveals that models without the application of motion energy may fail to capture the crucial characteristics of the sub-actions. For instance, in the "run + call" action, the model neglects the characteristic motion of making a phone call when motion energy is not used to emphasize the significant parts of the sub-action. This further demonstrates the critical role that motion energy plays in accurately capturing and preserving the distinctive characteristics of each sub-action in the final composed motion.